OpenStudy (anonymous):

Find the variation constant in which y varies inversely as x, and the following condition exists: y=20 when x=1/4

OpenStudy (amistre64):

y = k/x 20 = k/(1/4) 20 = k(4) 20/4 = k 5 = k